OpenStudy (anonymous):

Differentiate: y = (v^3 - 2v[sqrt[v]])/v

OpenStudy (anonymous):

\[\large y = \frac{v^3 -2v \sqrt{v}}{v}\] I would first simplify by factoring: \[\large y= \frac{\cancel v (v^2-2\sqrt{v})}{\cancel v}\] Then convert any roots to exponential form: \[\large y= v^2-2v^{\frac{1}{2}}\] Hopefully it is manageable from here

OpenStudy (swag):

nice work

OpenStudy (anonymous):

nice work..tengs :)